# Time arrow in open-boundary one-dimensional stochastic dynamics

### v2

@@ Abstract
We consider the finite-timestep Brownian dynamics of a single particle confined in one dimension, with a nonuniform temperature profile. In such an open-boundary scenario, one cannot observe any net probability current in the nonequilibrium steady state (NESS). On the other hand, the nonequilibrium nature of this system is exhibited through the asymmetry in forward and backward transition probabilities, as is reported in this work through the stochastic simulation analysis and theoretical arguments. The irreversibility becomes prominent nearby the temperature interface. We propose that the observed irreversibility can be accounted for via a virtual-gyration scenario, while the collapse of virtual gyrations upon the one-dimensional coordinate leads to the absence of probability current.
@@ I Introduction
In one-dimensional systems, persistent average gyration has been studied since the famous Büttiker-Landauer discussions in the 1980s[10, 11]. Through the closed-loop topology, a nonvanishing probability current can be achieved through a proper setting of temperature and potential profiles. The directed average dynamics, as an indicator of time irreversibility, is an important characteristic of the second law in thermodynamics, or its recent variants in stochastic dynamics, the fluctuation theorems[12, 13, 14, 15, 16]. In the Büttiker-Landauer engine, as well as in two-dimensional Brownian gyrators, it can be shown that the average net heat dissipation along the gyrating loop is positive. This is in agreement with the second law in the sense that the overall change in total entropy along a closed autonomous cycle is, on average, positive. Through the successful development of fluctuation theorems, scientists proceed to rephrase the second law of thermodynamics with more probability usage.
@@
If one considers the one-dimensional stochastic dynamics, just like the system of a Büttiker-Landauer engine, but with an open-boundary condition instead, the steady-state dynamics appears trivial with a zero average probability current over all positions. On the other hand, regarding thermodynamics, the system must stay in nonequilibrium under a nonuniform temperature profile. Without a persistent gyrating current, one may be tempted to search about the signature of time irreversibility for such a system. In this current work, we study the one-dimensional open-boundary system with discrete transition timesteps. We report that time irreversibility can indeed be observed, and virtual “gyrations” can occur on the local scale, for transitions near the temperature discontinuity.
@@
The nonequilibrium nature in similar temperature-gradient systems has been discussed in Refs. [8, 17, 18, 19]. These studies revealed the discrepancy in overdamped stochastic dynamics, as well as its shortage of contribution in total entropy dissipation of temperature-gradient systems. For example, in the work of Celani et al.[18], they obtained in underdamped stochastic dynamics an additional source of time irreversibility, or “entropy anomaly” in overall dissipation. This irreversible behavior is attributed to the circulating steady-state behavior in the phase space, and such an entropy anomaly persists even if one proceeds towards the overdamped limit. In our current work, we consider the overdamped stochastic dynamics only, and we find yet another source of time irreversibility, which can arise due to the finiteness of time intervals between successive thermal kicks. And in contrast to the work in [18], which assumes a smooth temperature variation over space, we choose to study the other extreme, in which temperature changes abruptly.
@@
Our work is structured as follows: in Sec. II, we introduce our system of study, its dynamic stochastic equation, as well as the corresponding Fokker-Planck equation analysis that works for infinitesimal timesteps. In Sec. III, we compute the finite-timestep transition probabilities and demonstrate through a crude approximation that time reversibility, or its synonym, detailed balance, may be violated. We go on to examine the violation of detailed balance through theoretical arguments. In Sec. IV we present our simulation result that confirms the existence of time irreversibility, and meanwhile, the steady-state probability distribution obtained from simulations exhibits deviation from the Fokker-Planck equation analysis. In Sec. V, we seek to fit the result of the probability distribution to that of a corresponding continuous-timestep model, through the introduction of an effective temperature profile. We remark that the smoothed-out characteristic of effective temperature can be interpreted by an averaging effect of the temperature profile over a single diffusion timestep.
@@ II System
We consider a Brownian particle moving in one dimension, its movement being dictated by the Langevin equation
@@
Thus, a Brownian particle of position $x$ is subject to a smooth potential $U(x)$, while the last term in the equation above gives rise to the random force. This fluctuating force is Gaussian white and uncorrelated, namely, $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$. For simplicity, all the physical quantities we use in this work are set to be dimensionless. Under this representation, the Boltzmann constant and $\gamma$ are both assigned to be unity. The latter can be achieved through a rescaling of time by the factor $1/\gamma$. The temperature is profiled by $T(x)=T_{H}$ for $x>x_{0}$ and $T(x)=T_{C}$ for $x\leq x_{0}$. Note that the product $\sqrt{T(x)}\,\xi$ in Eq. (1) adopts Itô’s scheme.
@@
represents the probability current. Upon reaching the steady state, $j=\rm{constant}=0$ everywhere in our one-dimensional open-boundary system. Therefore, within each constant-temperature domain, the probability density is proportional to the Boltzmann distribution. To determine the coefficients for the probability densities of these two regions, we consider an infinitesimal transition across the temperature discontinuity $x=x_{0}$. Since Eq. (3) implies $d(TP)+PdU=0$ in the NESS, if $PdU$ is negligible for infinitesimal transition step, so is $d(TP)=0$. Therefore, one has
@@
where $\beta\equiv 1/T$, and $U_{0}$ is the potential energy right at the temperature discontinuity.
Note that the FPE represents a special case of the master equation, following the truncation of the Kramers-Moyal expansion[20, 21] up to the second order. The legitimacy of truncation is based on the underlying assumption that transitions are local. In particular, for the one-dimensional case, the FPE can faithfully represent the statistics of random walks over nodes with a chain-like connecting structure as illustrated in Fig. 1(a). Regarding this open-chain structure, its steady-state distribution meets the detailed-balance criterion if the probability vanishes at far distances. However, in the following sections, we demonstrate that the detailed-balance condition, and hence time reversibility, fail in this model if random kicks occur over discrete timesteps and the temperature profile exhibits discontinuity. In such a case, the open-chain structure cannot represent the system faithfully. Meanwhile, this failure also leads to a notable deviation from the FPE analysis.
Figure 1: Illustrations of two scenarios. Arrows represent probability currents. Points $C$, $B$, $A$ represent successive discrete positions from left to right, while point $A$ is at a higher temperature. (a) The chain-connecting scenario. Note that the net steady state current between nodes has to be zero everywhere (detailed balance). (b) The loop scenario. In finite-timestep stochastic dynamics, point $A$ can reach $C$ directly over one single timestep $\Delta t$, thanks to the large diffusion parameter in $A$. This feature gives rise to the loop-like scenario and can contribute to a virtually gyrating current at the steady state.
@@ III Transition Probability; Violation of Detailed Balance
Now, let us study the Brownian motion again. This time, the random kicks emerge over discrete timesteps. Consider a transition from position $x$ to $x^{\prime}$ over timestep $\Delta t$. The Langevin equation then reads
@@
where $F=-\frac{dU}{dx}$ represents the conservative force, and $\xi\,\Delta t$ corresponds to the noise of a typical Wiener process. From Eq. (6), one derives the conditional transition probability
@@
To get a quick look at time irreversibility, let us assume for the time being that the solution of FPE is valid. With the use of Eq. (5) and (9), one then has
$\ln\frac{P(x^{\prime},x)}{P(x,x^{\prime})}$ $=\ln\left[\frac{P(x)}{P(x^{\prime})}\cdot\frac{P(x^{\prime}|x)}{P(x|x^{\prime})}\right]$
$\approx\frac{3}{2}\ln\frac{T^{\prime}}{T}+\frac{\Delta t}{4}\cdot\left(\frac{1}{T^{\prime}}-\frac{1}{T}\right)\cdot(\dot{x}^{2}+F^{2})-\frac{\Delta U}{2}\left(\frac{1}{T}+\frac{1}{T^{\prime}}\right)-\frac{1}{T}(U-U_{0})+\frac{1}{T^{\prime}}(U^{\prime}-U_{0})$
where $P(x^{\prime},x)$ represents the joint probability that the random walker is at site $x$ at time $t$ and $x^{\prime}$ at time $t+\Delta t$. Therefore, the detailed-balance condition holds for a uniform-temperature profile, at least to the order of magnitude considered here. On the other hand, Eq. (10) predicts time irreversibility for transitions across a temperature discontinuity.
Next, we proceed to show again the existence of time-irreversibility, or equivalently, detailed-balance violation, this time without assuming the validity of Eq. (5). Let us consider three neighboring sites $A,B$ and $C$, while the temperature at site $A$ is $T_{H}$ and the other two sites at a lower temperature $T_{C}$. The distances between $AB$ and $BC$ are identically equal, and we denote this distance to be $\Delta x$. Let us first consider the identity
$\frac{P(A,B)P(B,C)P(C,A)}{P(B,A)P(C,B)P(A,C)}=\frac{P(A|B)P(B|C)P(C|A)}{P(B|A)P(C|B)P(A|C)}\,.$ (11)
With Eq. (9), it leads to
@@
In our work we choose $(\Delta x)^{2}=O(\Delta t)$. Therefore, the magnitude of $\Delta x$ is equal to the diffusive scale over time $\Delta t$ if $O(T_{H})=O(T_{C})=1$. In this case, the first term in the final line of Eq. (12) outweighs the second term, and one observes significant violation in the detailed-balance condition: at least one forward-to-backward ratio of the $(A,B)$, $(B,C)$, $(C,A)$ pairs is not equal to 1.
Finally, we argue that the detailed-balance violation can occur for transitions over constant-temperature sites. Assuming that for our system of consideration, the detailed-balance condition still holds for constant-temperature transitions, then there would be no probability current between constant-temperature nodes. Let us expand our nodes of consideration such that the new node $A^{\prime}$ consists of all high-temperature sites, $B^{\prime}$ represents site $B$ only, while $C^{\prime}$ consists of all low-temperature sites other than site $B$. Thus, the assumption of detailed balance between $B^{\prime}$ and $C^{\prime}$ leads to the conclusion of no probability current between them. Yet a 3-node steady state cannot be achieved in this manner, if there exist nonzero currents between $(A^{\prime},B^{\prime})$ and/or $(A^{\prime},C^{\prime})$ (see Fig. 1(b)). Therefore,, we come to the conclusion that the detailed balance can be violated even between constant-temperature nodes. And in turn, it implies the failure of Eqs. (10) for constant-temperature transitions. This failure originates from the deviation in steady-state probability distribution from Eq. (5).
@@ IV.1 Joint Transition Probability
In this work, we choose $\delta=0.1$, which is similar to the diffusing distance over a single timestep $\Delta t$. The result, shown in Fig. 2, indicates significant deviation of $R(x)$ from unity near the temperature discontinuity, signifying the violation of detailed balance. Note that for single-step transitions over identical-temperature regions, Eq. (9) reduces to $-\Delta U/T$. According to this prediction, the ratio of joint probabilities $R(x)$ can be represented as
@@ IV.2 NESS probability distribution
[…]
@@ IV.3 Study in NESS probability current
The result of the computed $j(x)$, as presented in Fig. 4, exhibits noticeable deviation nearby the temperature discontinuity. The deficit is attributed to the fact that the computation scheme fails to count certain events. To be more specific, in the current model of description, the events that are generated from diffusions in the high-temperature region possess higher mobility. And these “fast” events, of which the random walker “bypasses” the monitoring bin over timestep $\Delta t$, often fail to be counted in the calculation of average velocity at the midpoints of their paths using Eq. (16) (see Fig. 1(b)). The deficit in the calculated probability current is thus more prominent nearby the temperature interface, where the calculated average velocity becomes positive.
@@ V Modified Temperature Profile and Probability Distribution
VI Summary
Through this study, we demonstrate the existence of time irreversibility in one-dimensional open-boundary stochastic dynamics when random thermal kicks occur over finite time intervals. Similar to the discussions in other areas of nonequilibrium steady-state stochastic dynamics, the irreversibility observed can be represented via some probability current “gyrating” among positions nearby the interface of temperature discontinuity (see Fig. 1(b)). The virtual gyration arises due to the mismatched diffusion distances between the high- and low-temperature adjacent regions. In terms of probability, the joint probabilities of forward and backward transitions exhibit differences nearby the temperature discontinuity, while the steady-state probability distribution in position also deviates from the description of the typical FPE analysis.
The characteristic width $w$, within which the aforementioned nonequilibrium phenomena could be observed, would shrink as the timestep between successive thermal kicks becomes smaller. From Eq. (18), one finds that $w^{2}/\Delta t$ bears the same order of magnitude as the diffusion coefficient. Taking the example of colloidal dispersions in water, this means that the time interval between successive kicks has to be in the order of seconds for $w$ to lie within the observation limit. While this is largely deviated from the natural condition for colloidal suspensions, the nonequilibrium facts discussed here may be observed more likely in systems with artificial noises[22, 23, 24, 25, 26, 27]. The realm of stochastic dynamics with discrete random kicks can be commonly observed in both natural and laboratory systems. Yet more extensive studies need to be conducted in order to explore the potentially distinctive behavior of its continuous-time counterpart. Finally, we remark that while we focus on the open-boundary one-dimensional system, the mechanism responsible for the irreversibility remains valid for the closed-loop Büttiker-Landauer engine, or higher-dimensional stochastic systems with temperature discontinuity.

### v3

@@ Abstract
We consider the finite-timestep Brownian dynamics of a single particle confined in one dimension, while the system has a nonuniform temperature profile. Under the open-boundary condition, one cannot observe any net probability current in the nonequilibrium steady state (NESS). On the other hand, the nonequilibrium nature of this system is revealed through the asymmetry in forward and backward transition probabilities, as is reported in this work through the simulation and theoretic analysis. In particular, our result shows prominent time irreversibility nearby the temperature interface. We propose a hidden-gyration scenario that demonstrates the source of irreversibility, while the collapse of gyrations on the one-dimensional coordinate accounts for the absence of probability current.
@@ I Introduction
In one-dimensional systems, persistent average gyration has been studied since the famous Büttiker-Landauer discussions in the 1980s[10, 11]. Through the closed-loop topology, a nonvanishing probability current can be achieved through a proper setting of temperature and potential profiles. The directed average dynamics, as an indicator of time irreversibility, is an important characteristic of the second law in thermodynamics, or its recent variants in stochastic dynamics, the fluctuation theorems[12, 13, 14, 15, 16]. In the Büttiker-Landauer engine, as well as in two-dimensional Brownian gyrators, it can be shown that the average net heat dissipation along the gyrating loop is positive. This is in agreement with the second law in the sense that the overall change in total entropy along a closed autonomous cycle is, on average, positive.
@@
In contrast to the examples above, the time arrow is seemingly absent regarding the stochastic dynamics in a one-dimensional open-boundary system. The steady-state dynamics in this system is trivial with a zero average probability current over all positions. On the other hand, the system cannot reach thermal equilibrium under a nonuniform temperature profile. Without a persistent gyrating current, one needs to search for an alternative signature of time irreversibility. In this current work, we study the one-dimensional open-boundary system with discrete transition timesteps. Note that the use of discrete thermal kicks in stochastic dynamics is ideal for the discussions of granular media, rarefied gases, and molecular-motor movements[17, 18, 19]. We conclude that time irreversibility can be observed in this platform, and gyrating behavior can be found at the local scale, for transitions near the temperature discontinuity.
@@
The nonequilibrium nature in similar temperature-gradient systems has been discussed in Refs. [8, 20, 21]. These studies revealed the discrepancy in overdamped stochastic dynamics, as well as its shortage of contribution in total entropy dissipation of temperature-gradient systems. For example, in the work of Celani et al.[20], they obtained in underdamped stochastic dynamics an additional source of time irreversibility, or “entropy anomaly” in overall dissipation. This irreversible behavior is attributed to the circulating steady-state behavior in the phase space, and such an entropy anomaly persists even if one proceeds towards the overdamped limit. In our current work, we consider the overdamped stochastic dynamics only, and we find yet another source of time irreversibility, which can arise due to the finiteness of time intervals between successive thermal kicks. And in contrast to the work in [20], which assumes a smooth temperature variation over space, our study is focused on the scenario where temperature changes discontinuously.
@@
Our work is structured as follows: first we introduce our system of study, its dynamic stochastic equation, as well as the corresponding Fokker-Planck equation analysis that works for infinitesimal timesteps. Then we compute the finite-timestep transition probabilities and demonstrate through a crude approximation that time reversibility, or its synonym, detailed balance, may be violated. We go on to examine the violation of detailed balance through theoretical arguments, and the simulation result also confirms the existence of time irreversibility. Meanwhile, the steady-state probability distribution obtained from the simulation shows deviation from the Fokker-Planck equation analysis. We then seek to fit the result of the probability distribution to that of a corresponding continuous-timestep model. This is achieved through the introduction of an effective temperature with a smoothed-out profile.
@@ II System
We consider a Brownian particle moving in one dimension. If the thermal kicks occur continuously, its movement can be depicted by the Langevin equation
@@
Thus, a Brownian particle of position $x$ is subject to a smooth potential $U(x)$, while the last term in the equation represents the source of the random force. This fluctuating force is Gaussian white and uncorrelated, namely, $\langle\xi(t)\xi(t^{\prime})\rangle=\delta(t-t^{\prime})$. For simplicity, all the physical quantities we use in this work are made dimensionless. Under this representation, the Boltzmann constant and $\gamma$ are both set to be unity. The latter can be achieved through a rescaling of time by the factor $1/\gamma$. The temperature is profiled by $T(x)=T_{H}$ for $x>x_{0}$ and $T(x)=T_{C}$ for $x\leq x_{0}$. Note that the product $\sqrt{T(x)}\,\xi$ in Eq. (1) adopts Itô’s scheme.
@@
represents the probability current. At the steady state, $j=\rm{constant}=0$ everywhere in the one-dimensional open-boundary system. Therefore, within each constant-temperature domain, the probability density is proportional to the Boltzmann distribution. To determine the prefactors at these two regions, we consider an infinitesimal transition across the temperature discontinuity $x=x_{0}$. Since Eq. (3) implies $d(TP)+PdU=0$ in the NESS, if $PdU$ is negligible for infinitesimal transition step, so is $d(TP)=0$. Therefore, one has
@@
where $\beta\equiv 1/T$, and $U_{0}$ is the potential energy right at the temperature discontinuity[22].
Note that the FPE represents a special case of the master equation, following the truncation of the Kramers-Moyal expansion[23, 24] up to the second order. The legitimacy of truncation is based on the underlying assumption that transitions are local. In particular, for 1-dim random walks, the FPE approach is valid for models with chain-like connecting structures (see Fig. 1(a)). Regarding an open-boundary system with a chain-like structure, its steady-state dynamics meets the detailed-balance criterion as long as the probability vanishes at far distances.
In the following sections, we study the 1-dim Brownian motion under discrete-timestep thermal kicks. We observe the failure of the detailed-balance condition, while the failure is most prominent for transitions across a temperature discontinuity. This failure is attributed to the fact that our discussed regime cannot be faithfully modeled by a open-chain structure. The absence of an open-chain structure can also account for a possible deviation from the FPE analysis.
Figure 1: Illustrations about the two scenarios. Arrows represent probability currents (width signifying magnitude). Points $C$, $B$, AA represent successive discrete positions from left to right, while point AA is at a higher temperature. (a) The chain-connecting scenario. Note that the net steady-state current between nodes has to be zero everywhere (detailed balance). (b) The loop scenario. In finite-timestep stochastic dynamics, thanks to the large diffusion parameter at AA, point AA can reach $C$ directly over one single timestep $\Delta t$. This feature gives rise to the loop-like scenario and contributes to a “hidden” gyrating current at the steady state.
@@ III Transition Probability; Violation of Detailed Balance
Now we study the stochastic dynamics in the regime where random kicks emerge over discrete timesteps. Consider a transition from position $x$ to $x^{\prime}$ over timestep $\Delta t$. The Langevin equation then reads
@@
where $F=-\frac{dU}{dx}$ represents the conservative force, and $\xi\,\Delta t$ corresponds to the noise in a typical Wiener process. From Eq. (6), one derives the conditional transition probability
@@
Next, we proceed to demonstrate the time irreversibility, or equivalently, detailed-balance violation. Let us consider three neighboring sites $A,B$ and $C$, while the temperature at site $A$ is $T_{H}$ and the other two sites at a lower temperature $T_{C}$. The distances between $AB$ and $BC$ are equal, and we denote this distance to be $\Delta x$. Let us first consider the identity
$\frac{P(A,B)P(B,C)P(C,A)}{P(B,A)P(C,B)P(A,C)}=\frac{P(A|B)P(B|C)P(C|A)}{P(B|A)P(C|B)P(A|C)}\,$, (10)
where $P(A,B)$ represents the joint probability that the random walker is at site $B$ at time $t$ and $A$ at a later time $t+\Delta t$. With Eq. (9), it leads to
@@
In our work we choose $(\Delta x)^{2}=O(\Delta t)$. Therefore, the order of magnitude in $\Delta x$ is comparable to the diffusive scale over time $\Delta t$. In this case, the first term in the final line of Eq. (11) outweighs the second term, and one observes significant violation in the detailed-balance condition: at least one forward-to-backward ratio among the $(A,B)$, $(B,C)$, $(C,A)$ pairs is not equal to one.
While the violation of detailed balance introduced above is attributed to the temperature change over sites, one may guess that the violation is present only for transitions straight across the temperature interface. We now proceed to argue that in fact, the detailed-balance violation can occur for transitions across constant-temperature sites. Let us temporarily assume that for our system of consideration, the detailed-balance condition still holds for all constant-temperature transitions. As a result, the probability current between constant-temperature neighboring sites must be zero. Let us enlarge our sites of consideration such that the “node” $A^{\prime}$ includes all high-temperature sites, node $B^{\prime}$ includes site $B$ only, and node $C^{\prime}$ includes all low-temperature sites other than site $B$. According to the assumption, there is be no probability current between nodes $B^{\prime}$ and $C^{\prime}$. However, this result is contradicting, since a three-way steady state cannot be achieved in this manner: if there exist some nonzero current between any of the two nodes, the same amount of current must circulate through the three nodes (see Fig. 1(b)). Therefore, we come to the conclusion that the detailed-balance condition must be violated between some constant-temperature nodes.
The three-way circulation in probability current described above is as a demonstration of the “hidden” gyration: the violation of detailed balance between successive sites suggests that the current can gyrate among local sites. On the other hand, the net probability current, i.e., the projection of current upon the one-dimensional coordinate, must be absent in the NESS dynamics of this open-boundary system.
Finally, we remark that if the solution of FPE (Eq. (5)) is valid, one then has
$\ln\frac{P(x^{\prime},x)}{P(x,x^{\prime})}$ $\approx\frac{3}{2}\ln\frac{T^{\prime}}{T}+\frac{\Delta t}{4}\cdot\left(\frac{1}{T^{\prime}}-\frac{1}{T}\right)\cdot(\dot{x}^{2}+F^{2})-\frac{\Delta U}{2}\left(\frac{1}{T}+\frac{1}{T^{\prime}}\right)-\frac{1}{T}(U-U_{0})+\frac{1}{T^{\prime}}(U^{\prime}-U_{0})$
which will vanish for transitions over constant-temperature sites. But this contradicts with our prediction in the previous paragraph. Thus one concludes that for finite-timestep stochastic dynamics, its steady-state probability distribution may deviate from Eq. (5), as will be shown in our simulation results.
@@ IV.1 Joint Transition Probability
In this work, we choose $\delta=0.1$, which is similar to the diffusing distance over a single timestep $\Delta t$. The result, shown in Fig. 2, indicates significant deviation of $R(x)$ from unity near the temperature discontinuity, signifying the violation of detailed balance. Note that for single-step transitions over identical-temperature regions, Eq. (9) reduces to $\Delta U/T$. According to this prediction, the ratio of joint probabilities $R(x)$ can be represented as
@@ IV.2 NESS probability distribution
$P(x)=\frac{1}{Z}\exp\left(-\int_{-\infty}^{x}\frac{\frac{d}{dx}(U(x)+T(x))}{T(x)}dx\right)\,$, (15)
@@ IV.3 Study in NESS probability current
The result of the computed $j(x)$, as presented in Fig. 4, exhibits noticeable deviation near the temperature discontinuity. The deficit is attributed to the fact that aforementioned computing scheme in $v$ and thus $j$ can miss certain events. To be more specific, in the current model of description, the events that are generated from diffusions in the high-temperature region possess higher mobility. And these “fast” events, of which the random walker “bypasses” the monitoring bin over timestep $\Delta t$, often fail to be counted in the calculation of average velocity at the midpoints of their paths using Eq. (16) (see Fig. 1(b)). The deficit in the calculated probability current is thus more prominent nearby the temperature interface, where the calculated average velocity becomes positive.
@@ V Modified Temperature Profile and Probability Distribution
VI Discussion and Summary
Throughout this study, we demonstrate the existence of time irreversibility in one-dimensional open-boundary stochastic dynamics when random thermal kicks occur over finite time intervals. Similar to the discussions in other areas of nonequilibrium steady-state stochastic dynamics, the irreversibility observed here can be represented via some “gyrating” probability current among positions near the temperature discontinuity (see Fig. 1(b)). The “hidden” gyration arises due to the mismatched diffusion distances between the high- and low-temperature adjacent regions. In terms of probability, the joint probabilities of forward and backward transitions exhibit differences near the temperature discontinuity, while the steady-state probability distribution in position also deviates from the description of the typical FPE analysis.
The characteristic width $w$, within which the aforementioned nonequilibrium phenomena could be observed, would shrink as the timestep between successive thermal kicks becomes smaller. From Eq. (18), one finds that $w^{2}/\Delta t$ bears the same order of magnitude as the diffusion coefficient. In colloidal suspensions, this source of time irreversibility is difficult to observe, because the thermal kicks occur almost continuously. On the other hand, the stochastic dynamics under discrete random kicks is commonly found in many natural and artificial systems, such as granular materials, seismic activities, or systems with artificial noises[18, 25, 26, 27, 28, 29, 30]. In these systems, it is more likely to observe the aforementioned time irreversibility.
Finally, we remark that while we focus on the open-boundary one-dimensional system, the mechanism responsible for the irreversibility remains valid for the closed-loop Büttiker-Landauer engine, or higher-dimensional stochastic systems with temperature discontinuity. Our current study may provide a guideline for future works towards a comprehensive understanding of NESS stochastic dynamics with discrete-timestep thermal noise, as well as its distinct behavior compared to the continuous-time counterpart.